# Physics-Constrained Fine-Tuning of Flow-Matching Models for Generation and Inverse Problems

### v1

@@ Appendix C Method: Implementation Details
In this section, we provide further details into the implementation of our method. This includes neural network architectures, the specific design of test functions for the computation of the weak residuals, numerical heuristics, and finally the full training algorithm for fine-tuning.
@@
All relevant code is implemented using Python 3.12.3, specifically, neural architectures are implemented in PyTorch (Paszke et al. 2019) version 2.7.
Inverse Predictor $\varphi$
We parametrise the inverse map $\varphi$ with a two-layer U-FNO, mirroring the spectral–spatial bias of the forward backbone. All other parameters are identical to the respective U-FNO for the vector field. For the 2D case the network operates on $(H\times W)$ tensors and outputs a permeability field of identical resolution. In the 3D setting we considered a time-independent parameter. Therefore, we collapse the temporal axis into input channels and apply a 2D U-FNO that performs spectral convolutions and U-Net operations solely in the spatial dimensions, yielding a $(H\times W)$ speed-of-sound map at substantially lower computational cost.
Architecture Modifications
Fine-tuning requires extending the original mapping
$v_{t,x}\leftarrow v(x,t)$
to a joint vector field
$(v_{t,x}^{\text{ft}},\,v_{t,\alpha}^{\text{ft}})\leftarrow v^{\text{ft}}(x,\alpha,t).$
Thus $\alpha$ must be supplied to the network as an additional input (so that $v_{t,x}^{\text{ft}}$ can be conditioned on it), and a second output channel $v_{t,\alpha}^{\text{ft}}$ has to be generated.
In practice a lightweight residual modification suffices. We first compute the padded U-FNO output as in the base model. A correction U-Net then receives three tensors—the preliminary vector field, the current state $x_{t}^{\text{ft}}$, and the parameter $\alpha_{t}^{\text{ft}}$—and produces an additive refinement. After removing the padding, the sum constitutes $v_{t,x}^{\text{ft}}$.
Learning $v_{t,\alpha}^{\text{ft}}$ from scratch would amount to training an entire denoising process for $\alpha$ during fine-tuning, which is impractical without supervised pre-training. Instead we adopt a residual strategy:
$v_{t,\alpha}^{\text{ft}}=v_{t,\alpha}^{\text{base}}+\mathrm{U\text{-}Net}(v_{t,\alpha}^{\text{base}},x_{t}^{\text{ft}},\alpha_{t}^{\text{ft}}).$
Both correctional U-Nets end with a $1\times 1$ convolution, such that pixel-wise corrections can be learned efficiently. Weights and biases of these final layers are initialised to zero so that, at the onset of fine-tuning, the fine-tuned model exactly reproduces the base model and departures are learned gradually.
Weak Residuals and Test Functions
@@
Let $u$ solve $\nabla\!\cdot\!\bigl{(}a\nabla u\bigr{)}-f=\mathcal{L}_{a}u=0$ on $\Omega\subset\mathbb{R}^{d}$ with homogeneous Dirichlet data $u|_{\partial\Omega}=0$. For any $\psi\in C^{1}_{0}(\Omega)$ we compute the $\text{L}^{2}$ inner product by multiplying with $\psi$ and integrating:
@@
where the boundary term vanishes because of $\psi|_{\partial\Omega}=0$. This expression only contains a first-order derivative of $u$ and can be used to compute in the computation of the weak residual.
Acoustic Wave Equation.
Let the pressure field
$p:\Omega\times[0,T]\to\mathbb{R}$
satisfy $\mathcal{L}_{c}p:=\partial_{t}^{2}p-c^{2}\Delta p=0$ with reflective (Neumann) boundary conditions $\partial_{n}p|_{\partial\Omega}=0$. For a test function $\psi\in C^{1}_{0}(\Omega\times[0,T])$ we form the $L^{2}$-inner product
$\langle\mathcal{L}_{c}p,\psi\rangle_{L^{2}(\Omega\times[0,T])}=\int_{0}^{T}\!\!\int_{\Omega}(\partial_{t}^{2}p-c^{2}\Delta p)\,\psi\;d\xi\,dt.$
Integrating the first term once in time and the second term once in space yields
$\langle\mathcal{L}_{c}p,\psi\rangle_{L^{2}}=$ $-\int_{0}^{T}\!\!\int_{\Omega}\partial_{t}p\,\partial_{t}\psi\;d\xi\,dt$
$+\int_{0}^{T}\!\!\int_{\Omega}c^{2}\,\nabla p\!\cdot\!\nabla\psi\;d\xi\,dt$
@@
$=$ $\int_{0}^{T}\!\!\int_{\Omega}\Bigl{(}-\partial_{t}p\,\partial_{t}\psi+c^{2}\,\nabla p\!\cdot\!\nabla\psi\Bigr{)}\,d\xi\,dt$
where all boundary terms vanish because $\psi$ is compactly supported in both space and time and $\partial_{n}p=0$ on $\partial\Omega$. Similar to the Darcy flow, the inner product now contains only first–order derivatives of $p$.
@@
where $r(x)=\sqrt{\sum_{j}(x_{j}-c_{j})^{2}/\sigma_{j}^{2}}$. The centre $c$ is uniform in $\Omega$ and axis length-scales $\sigma_{j}$ are uniform in $\sigma_{\min}\,\Delta_{j},\sigma_{\max}\,\Delta_{j}]$. By multiplying $\sigma_{\min}$ and $\sigma_{\max}$ with the grid spacing $\Delta_{j}$ of axis $j$, we obtain a parametrization that is intuitive to tune since the length-scales of the test functions are given in pixel units. $b\sim\mathrm{Ber}(p)$ randomly toggles a high-frequency wavelet factor that provides additional variability within the test functions and proved especially effective on noisy data.
@@
applied per axis. This legitimises the application of integration-by-parts in the derivation of the weak forms.
Note that in many systems the parameter itself scales the weak form. For Darcy flow, for instance, a perturbation that increases $\nabla u$ is linearly amplified by the permeability. If left unchecked, a joint optimiser can reduce the loss merely by driving the parameter toward smaller values. To avoid this degeneracy we normalise each test-function residual by the integral of $\alpha$ over the support of that function.
At training time we draw $N_{\text{test}}$ $\{\psi^{(i)}\}$ per residual evaluation and define the loss as
$\mathcal{R}_{\text{weak}}(x,\alpha)=\frac{1}{N_{\text{test}}}\sum_{i}\lvert\mathcal{\langle}\mathcal{L}_{\alpha}x,\psi^{(i)}\rangle_{L^{2}}\rvert^{2}.$
Heuristics
Similar to the heuristics stated in Domingo-Enrich et al. (2025), to allow for an efficient implementation, we deviate from the theoretical derivations outlined before. We lay them out in this section.

### v2

@@ Appendix D Method: Implementation Details
In this section, we provide further details into the implementation of our fine-tuning method. This includes neural network architectures, the specific design of test functions for the computation of the weak residuals, numerical heuristics, and finally the full training algorithm for fine-tuning.
@@
All relevant code is implemented using Python 3.12.3, specifically, neural architectures are implemented in PyTorch Paszke et al. (2019) version 2.7.
D.1 Inverse Predictor
We parametrise the inverse map $\varphi$ with a two-layer U-FNO, mirroring the spectral–spatial bias of the forward backbone. However, we increase the capacity of the U–Net components and use attention at the two lowest resolutions (stage indices 2 and 3). Since the inverse predictor only operates at $t=1$, we drop the temporal but keep the spatial conditioning. Exact parameters are stated in Table 5.
Table 5: U-FNO Inverse Predictor.
Hyperparameter Value
Number of layers 2
Padding mode reflect
Input channels 1 - 3
Time embedding dimension 0
Spatial embedding dimension 32
Output channels 1
Lifting width 256
Fourier modes $(k_{x},k_{y})$ [32, 32]
Spectral block width 32
U–Net base widths [64, 64, 96, 128]
U–Net embedding width 64
Attention stages [2, 3]
Attention heads 4
D.2 Architecture Modifications
Fine-tuning augments the base U–FNO map $x\mapsto v_{t,x}^{\text{base}}(x,t)$ to a joint, $\alpha$-conditioned vector field
$(v_{t,x}^{\text{ft}},\,v_{t,\alpha}^{\text{ft}})\;=\;v^{\text{ft}}(x,\alpha,t),$
implemented as residual corrections around the U–FNO core. Given the padded input $x_{\mathrm{pad}}$ (with time/space embeddings) and conditioning fields $\alpha$ and $v_{t,\alpha}^{\text{base}}$, the base feature stack $x_{\ast}$ is produced by the original spectral++skip++U–Net blocks. A first correction head (a U–Net) takes $x_{\ast},\,v_{t,x}^{\text{base}},\,\alpha]$ and outputs an additive refinement, yielding
$v_{t,x}^{\text{ft}}\;=\;v_{t,x}^{\text{base}}\;+\;\mathcal{U}_{x}\!\bigl(x_{\ast},\,v_{t,x}^{\text{base}},\,\alpha,\,t\bigr).$
After unpadding, a pixel-wise correction (lightweight channel-wise MLP) further adjusts $v_{t,x}^{\text{ft}}$ using local features and 2D positional channels,
$v_{t,x}^{\text{ft}}\;\leftarrow\;v_{t,x}^{\text{ft}}\;+\;\mathcal{M}_{x}\!\bigl(\text{pos}(x_{\ast},t),\,v_{t,x}^{\text{ft}}\bigr),$
which provides extra capacity for rapid adaptation without altering the global operator. For the parameter dynamics, we adopt a strictly residual strategy that conditions on both $\alpha$ and the baseline field:
$v_{t,\alpha}^{\text{ft}}\;=\;v_{t,\alpha}^{\text{base}}\;+\;\mathcal{U}_{\alpha}\!\bigl(x_{\mathrm{pad}},\,\alpha,\,v_{t,\alpha}^{\text{base}},\,t\bigr),$
where $\mathcal{U}_{\alpha}$ is a second U–Net. All correction heads are zero-initialized at their final projection layers, so that at the start of fine-tuning $v^{\text{ft}}\equiv(v_{t,x}^{\text{base}},\,v_{t,\alpha}^{\text{base}})$ and departures are learned smoothly. Table 6 lists the parameters of the correction U–Nets.
Table 6: Parameterization of the fine-tuning U–Net heads $\mathcal{U}_{x}$ and $\mathcal{U}_{\alpha}$.
Hyperparameter Value
U–Net base widths [64, 64, 96, 128]
Embedding width (time/aux) 64
Hidden lift (internal width) 256
Self-attention stages [2, 3]
Attention heads 4
Overall, the modifications to the base architecture add around 6M parameters to the model, resulting in a total $\approx$25M parameters.
D.3 Weak Residuals and Test Functions
@@
A pressure field $u$ that solves our Darcy flow equations fulfills $\nabla\!\cdot\!\bigl(a\nabla u\bigr)-f=\mathcal{L}_{a}u=0$} on $\Omega\subset\mathbb{R}^{d}$ with homogeneous Dirichlet data $u|_{\partial\Omega}=0$. For any $\psi\in C^{1}_{0}(\Omega)$ we compute the $\text{L}^{2}$ inner product by multiplying with $\psi$ and integrating:
@@
where the boundary term vanishes because of $\psi|_{\partial\Omega}=0$. This expression only contains a first-order derivative of $u$ and can be used to compute in the computation of the weak residual. To obtain a dimensionless, coefficient-scaled quantity comparable across locations, we normalize by the local mean permeability over the support of $\psi$,
$\overline{a}_{\psi}\;:=\;\frac{1}{|\operatorname{supp}\psi|}\int_{\operatorname{supp}\psi}a(\xi)\,d\xi,\qquad\widetilde{\mathcal{R}}[\psi]\;:=\;\frac{\mathcal{R}[\psi]}{\overline{a}_{\psi}}\,.$
In practice, $\operatorname{supp}\psi$ is the compact patch where $\psi$ (or its mollified variant) is nonzero, so that $\overline{a}_{\psi}$ captures the local coefficient scale used to normalize the residual.
Linear Elasticity.
For any compactly supported vector test $\psi\in C_{0}^{1}(\Omega;\mathbb{R}^{2})$, the weak residual is
@@
$\mathcal{R}[\psi]\;:=\;\langle\mathcal{L}_{E}u,\psi\rangle_{\mathrm{L}^{2}(\Omega)}\;=\;\int_{\Omega}\bigl(-\nabla\!\cdot\sigma(u;E,\nu)\bigr)\!\cdot\!\psi\,d\xi.$
A single integration by parts yields
$\mathcal{R}[\psi]=-\int_{\partial\Omega}\bigl(\sigma n\bigr)\!\cdot\!\psi\,dS\;+\;\int_{\Omega}\sigma:\nabla\psi\,d\xi\;=\;\int_{\Omega}\sigma:\nabla\psi\,d\xi,$
since $\psi|_{\partial\Omega}=0$. Here $\sigma:\nabla\psi$” is the Frobenius product and we denote the stress components by $\sigma_{xx},\sigma_{xy}(=\sigma_{yx}),\sigma_{yy}$. To reuse the same scalar test generator for both momentum equations, we restrict to tests that share a single scalar profile in both components. Concretely, we take
$\psi^{(x)}=(\psi,0),\qquad\psi^{(y)}=(0,\psi),\qquad\psi\in C_{0}^{1}(\Omega).$
With this restriction, residuals can be computed component-wise
$\langle\mathcal{L}_{E}u,\psi^{(x)}\rangle=\int_{\Omega}\bigl(\sigma_{xx}\,\partial_{\xi_{1}}\psi+\sigma_{xy}\,\partial_{\xi_{2}}\psi\bigr)\,d\xi,\qquad\langle\mathcal{L}_{E}u,\psi^{(y)}\rangle=\int_{\Omega}\bigl(\sigma_{xy}\,\partial_{\xi_{1}}\psi+\sigma_{yy}\,\partial_{\xi_{2}}\psi\bigr)\,d\xi.$
To obtain a dimensionless, coefficient-scaled quantity, we normalize by the local mean modulus over the support of $\psi$,
$\overline{E}_{\psi}:=\frac{1}{|\operatorname{supp}\psi|}\int_{\operatorname{supp}\psi}E(\xi)\,d\xi,\qquad\widetilde{\mathcal{R}}^{(x)}[\psi]:=\frac{\langle\mathcal{L}_{E}u,\psi^{(x)}\rangle}{\overline{E}_{\psi}},\quad\widetilde{\mathcal{R}}^{(y)}[\psi]:=\frac{\langle\mathcal{L}_{E}u,\psi^{(y)}\rangle}{\overline{E}_{\psi}}.$
For a family of tests $\{\psi_{k}\}_{k=1}^{N}$, the scalar residual used in experiments is the $\ell_{2}$ aggregation over components followed by averaging over tests. Therefore, the weak residual for the elasticity experiment is
$\mathcal{R}_{\mathrm{weak}}:=\frac{1}{N}\sum_{k=1}^{N}\Bigl(\bigl(\widetilde{\mathcal{R}}^{(x)}[\psi_{k}]\bigr)^{2}+\bigl(\widetilde{\mathcal{R}}^{(y)}[\psi_{k}]\bigr)^{2}\Bigr).$
Helmholtz.
For the Helmholtz experiments we consider complex-valued time–harmonic pressures $u:\Omega\to\mathbb{C}$ satisfying
$\mathcal{L}_{c}u\;:=\;-\Delta u(\xi)-\bigl(1-i\tan\delta\bigr)\,\kappa(\xi)^{2}u(\xi)-s(\xi)\;=\;0,\qquad\kappa(\xi)=\frac{\omega}{c(\xi)}\,$,
where $c(\xi)$ is the sound speed, $\omega>0$ the angular frequency, $\tan\delta\geq 0$ the loss tangent, and $s(\xi)$ a fixed source. Writing $u=u_{R}+iu_{I}$ with real and imaginary parts $u_{R},u_{I}:\Omega\to\mathbb{R}$, we split the operator into its real and imaginary components
$\mathcal{L}_{c}u=\bigl(\mathcal{L}_{c}^{(R)}u_{R}+\mathcal{C}^{(R)}u_{I}-s\bigr)+i\bigl(\mathcal{L}_{c}^{(I)}u_{I}+\mathcal{C}^{(I)}u_{R}\bigr),$
with
$\mathcal{L}_{c}^{(R)}u_{R}:=-\Delta u_{R}-\kappa^{2}u_{R},\qquad\mathcal{L}_{c}^{(I)}u_{I}:=-\Delta u_{I}-\kappa^{2}u_{I},$
$\mathcal{C}^{(R)}u_{I}:=\alpha_{I}u_{I},\qquad\mathcal{C}^{(I)}u_{R}:=-\alpha_{I}u_{R},$
where $\alpha_{R}:=\kappa^{2}$ and $\alpha_{I}:=-\tan\delta\,\kappa^{2}$. For any scalar test function $\psi\in C_{0}^{1}(\Omega)$, we define real and imaginary weak residuals as
$\mathcal{R}^{(R)}[\psi]:=\bigl\langle\mathcal{L}_{c}^{(R)}u_{R}+\mathcal{C}^{(R)}u_{I}-s,\;\psi\bigr\rangle_{L^{2}(\Omega)},\qquad\mathcal{R}^{(I)}[\psi]:=\bigl\langle\mathcal{L}_{c}^{(I)}u_{I}+\mathcal{C}^{(I)}u_{R},\;\psi\bigr\rangle_{L^{2}(\Omega)}.$
A single integration by parts in the Laplacian terms gives
$\mathcal{R}^{(R)}[\psi]$ $=-\int_{\partial\Omega}(\nabla u_{R}\cdot n)\,\psi\,dS+\int_{\Omega}\bigl(\nabla u_{R}\cdot\nabla\psi-\alpha_{R}u_{R}\psi+\alpha_{I}u_{I}\psi-s\,\psi\bigr)\,d\xi,$
$\mathcal{R}^{(I)}[\psi]$ $=-\int_{\partial\Omega}(\nabla u_{I}\cdot n)\,\psi\,dS+\int_{\Omega}\bigl(\nabla u_{I}\cdot\nabla\psi-\alpha_{R}u_{I}\psi-\alpha_{I}u_{R}\psi\bigr)\,d\xi.$
Since $\psi$ is compactly supported in the interior, the boundary contributions vanish and we use only the volume terms in the residual. As in the elasticity case, we reuse the same scalar profile $\psi$ for both channels (real and imaginary parts) and evaluate $\mathcal{R}^{(R)}[\psi]$ and $\mathcal{R}^{(I)}[\psi]$ on identical compact patches to increase computational efficiency.
To obtain a dimensionless quantity that is comparable across locations and coefficient scales, we normalize by a local Helmholtz energy associated with the bilinear form
$E_{\psi}(u)\;:=\;\int_{\operatorname{supp}\psi}\Bigl(|\nabla u_{R}|^{2}+|\nabla u_{I}|^{2}+\kappa(\xi)^{2}\bigl(u_{R}(\xi)^{2}+u_{I}(\xi)^{2}\bigr)\Bigr)\,d\xi,$
which combines gradient and reaction contributions of both channels. The normalized residuals are
$\widetilde{\mathcal{R}}^{(R)}[\psi]:=\frac{\mathcal{R}^{(R)}[\psi]}{\sqrt{E_{\psi}(u)}},\qquad\widetilde{\mathcal{R}}^{(I)}[\psi]:=\frac{\mathcal{R}^{(I)}[\psi]}{\sqrt{E_{\psi}(u)}}.$
Given a family of scalar tests $\{\psi_{k}\}_{k=1}^{N}$, the scalar weak residual used in experiments is obtained by aggregating real and imaginary components and averaging over tests,
$\mathcal{R}_{\mathrm{weak}}:=\frac{1}{N}\sum_{k=1}^{N}\Bigl(\bigl(\widetilde{\mathcal{R}}^{(R)}[\psi_{k}]\bigr)^{2}+\bigl(\widetilde{\mathcal{R}}^{(I)}[\psi_{k}]\bigr)^{2}\Bigr).$
Stokes lid-driven cavity.
For the Stokes experiments, each sample contains a velocity–pressure pair $(u,p)$ solving the variable–viscosity incompressible Stokes system
$\mathcal{L}_{\nu}(u,p):=\bigl(-\nabla\!\cdot(\nu\nabla u)+\nabla p,\;\nabla\!\cdot u\bigr)=(0,0)\qquad\text{in }\Omega,$
with $\nu:\Omega\to\mathbb{R}_{>0}$ denoting the viscosity field. Writing $u=(u_{x},u_{y})$, the operator decomposes into
$\mathcal{L}_{\nu}^{(x)}(u,p)=-\nabla\!\cdot(\nu\nabla u_{x})+\partial_{\xi_{1}}p,\qquad\mathcal{L}_{\nu}^{(y)}(u,p)=-\nabla\!\cdot(\nu\nabla u_{y})+\partial_{\xi_{2}}p,$
$\mathcal{L}_{\nu}^{(\mathrm{div})}(u,p)=\partial_{\xi_{1}}u_{x}+\partial_{\xi_{2}}u_{y}.$
For any scalar test function $\psi\in C_{0}^{1}(\Omega)$, we form momentum and incompressibility residuals by testing each of these three equations with the same scalar profile,
$\mathcal{R}^{(x)}[\psi]:=\;\bigl\langle\mathcal{L}_{\nu}^{(x)}(u,p),\psi\bigr\rangle_{L^{2}(\Omega)},\qquad\mathcal{R}^{(y)}[\psi]:=\;\bigl\langle\mathcal{L}_{\nu}^{(y)}(u,p),\psi\bigr\rangle_{L^{2}(\Omega)},$
$\mathcal{R}^{(\mathrm{div})}[\psi]:=\;\bigl\langle\mathcal{L}_{\nu}^{(\mathrm{div})}(u,p),\psi\bigr\rangle_{L^{2}(\Omega)}.$
As in the Darcy and elasticity settings, one integration by parts isolates only first derivatives of the solution inside the volume integral. Since $\psi$ has compact support inside $\Omega$, all boundary terms vanish and we obtain the local weak forms
$\mathcal{R}^{(x)}[\psi]$ $=\int_{\Omega}\Bigl(\nu\,\nabla u_{x}\cdot\nabla\psi+(\partial_{\xi_{1}}p)\,\psi\Bigr)\,d\xi,$
$\mathcal{R}^{(y)}[\psi]$ $=\int_{\Omega}\Bigl(\nu\,\nabla u_{y}\cdot\nabla\psi+(\partial_{\xi_{2}}p)\,\psi\Bigr)\,d\xi,$
$\mathcal{R}^{(\mathrm{div})}[\psi]$ $=\int_{\Omega}\bigl(\partial_{\xi_{1}}u_{x}+\partial_{\xi_{2}}u_{y}\bigr)\,\psi\,d\xi.$
To obtain a dimensionless, coefficient–scaled residual comparable across locations, we normalize by the local viscosity–weighted kinetic energy over the support of $\psi$,
$E_{\psi}(u):=\int_{\operatorname{supp}\psi}\Bigl(\nu(\xi)\,\|\nabla u(\xi)\|_{F}^{2}+\|u(\xi)\|_{2}^{2}\Bigr)\,d\xi,$
where $\|\nabla u\|_{F}^{2}=|\partial_{\xi_{1}}u_{x}|^{2}+|\partial_{\xi_{2}}u_{x}|^{2}+|\partial_{\xi_{1}}u_{y}|^{2}+|\partial_{\xi_{2}}u_{y}|^{2}$. The normalized weak residuals are
$\widetilde{\mathcal{R}}^{(x)}[\psi]:=\frac{\mathcal{R}^{(x)}[\psi]}{\sqrt{E_{\psi}(u)}},\qquad\widetilde{\mathcal{R}}^{(y)}[\psi]:=\frac{\mathcal{R}^{(y)}[\psi]}{\sqrt{E_{\psi}(u)}},\qquad\widetilde{\mathcal{R}}^{(\mathrm{div})}[\psi]:=\frac{\mathcal{R}^{(\mathrm{div})}[\psi]}{\sqrt{E_{\psi}(u)}}.$
Finally, for a collection of scalar tests $\{\psi_{k}\}_{k=1}^{N}$, the scalar weak residual used in experiments aggregates momentum and incompressibility contributions and averages across tests:
$\mathcal{R}_{\mathrm{weak}}:=\frac{1}{N}\sum_{k=1}^{N}\Bigl(\bigl(\widetilde{\mathcal{R}}^{(x)}[\psi_{k}]\bigr)^{2}+\bigl(\widetilde{\mathcal{R}}^{(y)}[\psi_{k}]\bigr)^{2}+\bigl(\widetilde{\mathcal{R}}^{(\mathrm{div})}[\psi_{k}]\bigr)^{2}\Bigr).$
@@
where $r(x)=\sqrt{\sum_{j}(x_{j}-c_{j})^{2}/\sigma_{j}^{2}}$. Length-scales $\sigma_{j}$ are uniformly sampled from the range $\sigma_{\min}\,\Delta_{j},\sigma_{\max}\,\Delta_{j}]$. By multiplying $\sigma_{\min}$ and $\sigma_{\max}$ with the grid spacing $\Delta_{j}$ of axis $j$, we obtain a parametrization that is intuitive to tune since the length-scales of the test functions are given in pixel units. Instead of also sampling center points $c$ uniformly and independently from the full domain $\Omega$, we instead start from the grid coordinates of the data, considering each grid point as one center. We found that this way of ensuring spatial coverage improves training efficiency. To still retain stochasticity, we apply i.i.d jitter to each center point. $b\sim\mathrm{Ber}(p)$ randomly toggles a high-frequency wavelet factor that provides additional variability within the test functions and proved especially effective on noisy data.
@@
applied per axis. This legitimises the application of integration-by-parts in the derivation of the weak forms. At training time we sample a set of test functions $\{\psi^{(i)}\}_{i=1}^{N_{\text{test}}}$ independently per residual evaluation and define the loss as
$\mathcal{\mathcal{}}{R}_{\text{weak}}(x,\alpha)=\frac{1}{N_{\text{test}}}\sum_{i}\lvert\mathcal{\langle}\mathcal{L}_{\alpha}x,\psi^{(i)}\rangle_{L^{2}}\rvert^{2}.$
D.4 Adjoint Matching Details
As mentioned in the main paper, we introduce a scaling coefficient $\kappa$ that allows us to attenuate the magnitude of the noise variance. In this section, we provide a justification for using $\kappa>0$ and lay out numerical heuristics used in the implementation.